OpenStudy (anonymous):

The following ratios form a proportions 3/9 : 5/15 true or false

OpenStudy (anonymous):

what is 3/9 in simplest form?

OpenStudy (anonymous):

3/9 : 5/15 3/9=1/3 and 5/15=1/3 so, it is true